### 04 Mar 2026 v1

# A New Class of Geometric Analog Error Correction Codes for Crossbar Based In-Memory Computing

Abstract
Analog error correction codes have been proposed for analog in-memory computing on resistive crossbars, which can accelerate vector–matrix multiplication for machine learning. Unlike traditional communication or storage channels, this setting involves a mixed noise model with small perturbations and outlier errors. A number of analog codes have been proposed for handling a single outlier, and several constructions have also been developed to address multiple outliers. However, the set of available code families remains limited, covering only a narrow range of code lengths and dimensions. In this paper, we study a recently proposed family of geometric codes capable of handling multiple outliers, and develop a geometric analysis that characterizes their $m$-height profiles.
I Introduction
Analog in-memory computing is a cutting-edge technology that integrates data storage and computation directly within memory cells, enabling significant acceleration of deep neural network (DNN) computations [1, 2]. The primary motivation for analog in-memory computing is to overcome the "von Neumann bottleneck" by avoiding the need for massive data transfers between processors and memory [3, 4, 5]. This approach promises substantial improvements in speed and energy efficiency by exploiting the vector–matrix multiplications within crossbar architectures [6], a fundamental operation in DNNs.
Analog ECCs have been proposed to address the challenge of computing against errors [7, 8]. These codes are specifically designed to handle errors in codewords transmitted through channels that introduce two primary types of additive noise: limited-magnitude errors (LMEs) and unlimited-magnitude errors (UMEs). LMEs are small but widespread, arising from effects like cell-programming noise and random read/write disturbances in non-volatile memory (NVM) arrays. In contrast, UMEs, such as stuck cells or short cells, occur less frequently but can be significantly more disruptive. While DNNs can often tolerate minor, distributed noise, they are particularly vulnerable to large, isolated errors, making robust error correction essential for reliable in-memory computation [1].
Some existing analog ECC designs primarily target detection of a single UME [7, 8], while others address multiple UMEs [9]. In addition, some promising short codes that handle multiple UMEs have been obtained via genetic algorithms [10]. More recently, [11] proposed a new class of geometric codes including polygonal codes, dual polygonal codes, polyhedral codes, and dual polyhedral codes together with a linear-programming analysis of their error-handling capability. In this paper, we develop an alternative geometric analysis for the dual polygonal codes and dual polyhedral codes, and characterize their $m$-height profiles, recovering the same results as in [11]. Extending this geometric approach to polygonal codes and polyhedral codes remains an open problem.
II Preliminaries
Let $\mathcal{C}\subset\mathbb{R}^{n}$ be a real-valued linear code of dimension $k$ generated by a real $k\times n$ matrix $G$. Each codeword has the form $\mathbf{c}=\mathbf{u}G$, where $\mathbf{u}\in\mathbb{R}^{k}$ is an arbitrary input vector and $\mathbf{c}\in\mathbb{R}^{n}$. For a codeword $\mathbf{c}=(c_{0},c_{1},\ldots,c_{n-1})\in\mathcal{C}$, consider the magnitudes $|c_{0}|,|c_{1}|,\ldots,|c_{n-1}|$. Let $c_{(0)}\geq c_{(1)}\geq\cdots\geq c_{(n-1)}$ denote the elements of $\{|c_{j}|\}_{j=0}^{n-1}$ sorted in nonincreasing order.
The $m$-height of a codeword $\mathbf{c}$ is defined as the ratio of its largest-magnitude entry to its $(m+1)$-th largest-magnitude entry, i.e., $h_{m}(\mathbf{c})=\frac{c_{(0)}}{c_{(m)}}$. The $m$-height of the code $\mathcal{C}$ is then
$h_{m}(\mathcal{C})=\max_{\mathbf{c}\in\mathcal{C}}h_{m}(\mathbf{c}),$
namely the largest $m$-height attained over all codewords in $\mathcal{C}$. Theorem 2 of [10] gives a general method to compute $h_{m}(\mathcal{C})$ by solving a family of linear programs, and a simplified version of this approach is provided in [11].
Let $\delta$ and $\Delta$ be positive real thresholds with $\Delta>\delta>0$. An error vector $\bm{\varepsilon}=(\varepsilon_{0},\varepsilon_{1},\ldots,\varepsilon_{n-1})\in\mathbb{R}^{n}$ is called a LME if $\varepsilon_{i}\in[-\delta,\delta]$ for all $i\in\{0,1,\ldots,n-1\}$. Values outside the interval $\Delta,\Delta]$ are referred to as UMEs. The error-handling capability of a code $\mathcal{C}$ is characterized by the following parameters: (i) the number $\tau$ of outliers that can be located, (ii) the number $\tau+\sigma$ of outliers that can be detected, and (iii) the ratio $\Delta/\delta$. In terms of the height profile, the following necessary and sufficient condition relates these parameters [7]:
$\frac{\Delta}{\delta}\geq 2\bigl(h_{2\tau+\sigma}(\mathcal{C})+1\bigr).$
Thus, achieving smaller $m$-height directly translates to stronger UME location and detection capability.
III Dual Polygonal Codes
In this section, we construct a class of $k=2$ analog codes whose generator matrices consist of evenly spaced unit vectors over a half circle, which is the same as the code in Example 4 of [11].
Let $k=2$ and let $G=[{\bm{g}}_{0},\dots,{\bm{g}}_{n-1}]\in\mathbb{R}^{2\times n}$ consist of unit columns
${\bm{g}}_{j}=\begin{bmatrix}\cos\theta_{j}\\[2.0pt] \sin\theta_{j}\end{bmatrix},\qquad\theta_{j}\in[0,\pi).$
Choose $\theta_{j}=\frac{\pi}{n}\,j$, $j=0,1,\dots,n-1$. Then the set $\{{\bm{g}}_{j}\}$ forms an evenly spaced set of vectors over a half circle. An example with $n=3$ is illustrated in Fig. 1. The dashed lines mark the antipodal directions of the generator vectors.
${\bm{g}}_{0}$${\bm{g}}_{1}$${\bm{g}}_{2}$
Figure 1: An example of dual polygonal codes for $n=3$.
For a unit information vector ${\bm{u}}(\alpha)=(\cos\alpha,\sin\alpha)^{\top}$, the resulting codeword is ${\bm{c}}(\alpha)={\bm{u}}(\alpha)^{\top}G\in\mathbb{R}^{n}$, whose entries are $c_{j}(\alpha)=\cos(\theta_{j}-\alpha)$. Let $c_{(0)}(\alpha)\geq c_{(1)}(\alpha)\geq\cdots\geq c_{(n-1)}(\alpha)$ denote the elements of $\{|c_{j}(\alpha)|\}_{j=0}^{n-1}$ sorted in nonincreasing order. Then the $m$-height of the code is
$h_{m}(\mathcal{C})=\sup_{\alpha\in[0,2\pi)}\frac{c_{(0)}(\alpha)}{c_{(m)}(\alpha)}.$
Information directions $\alpha\in\bigl[0,\tfrac{\pi}{2n}\bigr]$ generate a subset of codewords. Due to symmetry, all other codewords can be obtained from this subset by suitable permutations and sign changes of the coordinates, which do not affect the $m$-height. As a result, the interval $\alpha\in\bigl[0,\tfrac{\pi}{2n}\bigr]$ constitutes a complete and sufficient domain for analyzing the $m$-height of the code, as illustrated in Fig. 1. Hence, the $m$-height simplifies to
$h_{m}(\mathcal{C})=\sup_{\alpha\in[0,2\pi)}\frac{c_{(0)}(\alpha)}{c_{(m)}(\alpha)}=\sup_{\alpha\in[0,\frac{\pi}{2n}]}\frac{c_{(0)}(\alpha)}{c_{(m)}(\alpha)}.$
Lemma III.1.
Let $n\geq 2$ and fix $\alpha\in\bigl[0,\tfrac{\pi}{2n}\bigr]$. Then the order statistics $c_{(k)}(\alpha)$ of $\lvert c_{j}(\alpha)\rvert$ are attained at the indices
$c_{(k)}(\alpha)=|c_{j_{k}}(\alpha)|,\qquad k=0,1,\ldots,n-1,$
where
$j_{k}=\begin{cases}0,&k=0,\\[4.0pt] \dfrac{k+1}{2},&\text{$k$ odd},\\[6.0pt] n-\dfrac{k}{2},&\text{$k$ even and }k\geq 2.\end{cases}$
Proof.
Recall that
$c_{j}(\alpha)=\cos\!\Bigl(\tfrac{\pi j}{n}-\alpha\Bigr),\qquad j=0,1,\ldots,n-1,$
and by definition $c_{(k)}(\alpha)$ denotes the $(k+1)$-th largest value among $\bigl\{\lvert c_{j}(\alpha)\rvert\bigr\}_{j=0}^{n-1}$.
Since $|\cos x|$ is $\pi$-periodic and even, it is convenient to reduce to the principal strip $(-\tfrac{\pi}{2},\tfrac{\pi}{2}]$. Define
$P(x)\;:=\;x-\pi\Big\lfloor\frac{x+\tfrac{\pi}{2}}{\pi}\Big\rfloor\in\bigl(-\tfrac{\pi}{2},\tfrac{\pi}{2}\bigr],$
so that
$\bigl|\cos P(x)\bigr|=\bigl|\cos x\bigr|\quad\text{for all }x\in\mathbb{R}.$
Let
$\mathcal{M}$ $=\bigl\{-\alpha+j\tfrac{\pi}{n}:j=0,1,\ldots,n-1\bigr\},$
$\mathcal{L}$ $=P(\mathcal{M})=\bigl\{\,P(-\alpha+j\tfrac{\pi}{n}):j=0,1,\ldots,n-1\,\bigr\}.$
Then
$\bigl\{\lvert c_{j}(\alpha)\rvert\bigr\}_{j=0}^{n-1}=\bigl\{\lvert\cos x\rvert:x\in\mathcal{M}\bigr\}$
$=\bigl\{\lvert\cos x\rvert:x\in\mathcal{L}\bigr\}=\bigl\{\cos x:x\in\mathcal{L}\bigr\},$
because $\mathcal{L}\subset(-\tfrac{\pi}{2},\tfrac{\pi}{2}]$ and $\cos x>0$ on this interval.
On $0,\tfrac{\pi}{2}]$ the function $x\mapsto\cos x$ is strictly decreasing, and $\cos x$ is even. Hence for $|\alpha_{1}|,|\alpha_{2}|\leq\tfrac{\pi}{2}$ we have
$\cos\alpha_{1}\geq\cos\alpha_{2}\iff|\alpha_{1}|\leq|\alpha_{2}|.$
Therefore, within $\mathcal{L}$, ordering the values $\{\cos x:x\in\mathcal{L}\}$ in nonincreasing order is equivalent to ordering the absolute values $\{|x|:x\in\mathcal{L}\}$ in nondecreasing order.
We now describe $\mathcal{L}$ explicitly. First, for $j=0,1,\ldots,\lfloor\tfrac{n}{2}\rfloor$ we have
$\alpha+j\tfrac{\pi}{n}\in\bigl(-\tfrac{\pi}{2},\tfrac{\pi}{2}\bigr],$
$P\!\Bigl(-\alpha+j\tfrac{\pi}{n}\Bigr)=-\alpha+j\tfrac{\pi}{n}.$
These contribute the points
$\alpha,\ \tfrac{\pi}{n}-\alpha,\ \tfrac{2\pi}{n}-\alpha,\ \tfrac{3\pi}{n}-\alpha,\ \ldots$
up to the last $j$ for which the expression stays in $(-\tfrac{\pi}{2},\tfrac{\pi}{2}]$.
Next, consider $j=\tfrac{n}{2}+r$ (if $n$ is even) or $j=\tfrac{n-1}{2}+1+r$ (if $n$ is odd); in either case we can write
$j=n-r,\qquad r=1,2,\ldots,\Bigl\lfloor\tfrac{n-1}{2}\Bigr\rfloor.$
Then
$\alpha+j\tfrac{\pi}{n}=-\alpha+(n-r)\tfrac{\pi}{n}=\pi-\Bigl(\tfrac{\pi r}{n}+\alpha\Bigr),$
which lies in $(\tfrac{\pi}{2},\pi)$ for the relevant $r$. Applying $P$ gives
$P\!\Bigl(-\alpha+(n-r)\tfrac{\pi}{n}\Bigr)=-\Bigl(\tfrac{\pi r}{n}+\alpha\Bigr),$
so these contribute the points
$\Bigl(\tfrac{\pi}{n}+\alpha\Bigr),\ -\Bigl(\tfrac{2\pi}{n}+\alpha\Bigr),\ -\Bigl(\tfrac{3\pi}{n}+\alpha\Bigr),\ \ldots.$
Collecting both parts, the elements of $\mathcal{L}$ in $(-\tfrac{\pi}{2},\tfrac{\pi}{2}]$ can be written as
$\alpha,\ \tfrac{\pi}{n}-\alpha,\ \tfrac{2\pi}{n}-\alpha,\ \ldots\quad\text{and}\quad-\Bigl(\tfrac{\pi}{n}+\alpha\Bigr),\ -\Bigl(\tfrac{2\pi}{n}+\alpha\Bigr),\ \ldots$
hence their absolute values form the interleaving sequence
$\alpha,\ \tfrac{\pi}{n}-\alpha,\ \tfrac{\pi}{n}+\alpha,\ \tfrac{2\pi}{n}-\alpha,\ \tfrac{2\pi}{n}+\alpha,\ \tfrac{3\pi}{n}-\alpha,\ \tfrac{3\pi}{n}+\alpha,\ \ldots,$
because for every integer $r\geq 1$ and every $\alpha\in[0,\tfrac{\pi}{2n}]$,
$r\tfrac{\pi}{n}-\alpha\;\leq\;r\tfrac{\pi}{n}+\alpha\;\leq\;(r+1)\tfrac{\pi}{n}-\alpha.$
We now map these distances back to indices $j$. By construction,
  • The smallest distance $\alpha$ corresponds to $j=0$.
  • The next distance $\tfrac{\pi}{n}-\alpha$ corresponds to $j=1$.
  • The distance $\tfrac{\pi}{n}+\alpha$ corresponds to j=n1j=n-1.
  • The distance $\tfrac{2\pi}{n}-\alpha$ corresponds to $j=2$.
  • The distance $\tfrac{2\pi}{n}+\alpha$ corresponds to j=n2j=n-2.
and so on. In general, the interleaving pattern of distances yields the index sequence
0, 1,n1, 2,n2, 3,n3,,0,\ 1,\ n-1,\ 2,\ n-2,\ 3,\ n-3,\ $\ldots$,
i.e.,
$j_{k}=\begin{cases}0,&k=0,\\[4.0pt] \dfrac{k+1}{2},&\text{$k$ odd},\\[6.0pt] n-\dfrac{k}{2},&\text{$k$ even and }k\geq 2.\end{cases}$
Since $c_{(k)}(\alpha)$ is, by definition, the $(k+1)$-th largest element of $\{\lvert c_{j}(\alpha)\rvert\}_{j=0}^{n-1}$, and we have just shown that the corresponding indices appear in the order $j_{0},j_{1},j_{2},\ldots$ given above, which proves the lemma. ∎
Thus, by Lemma III.1, the $m$-height computation can be simplified as follows.
$h_{m}(\mathcal{C})=\sup_{\alpha\in[0,\frac{\pi}{2n}]}\frac{c_{(0)}(\alpha)}{c_{(m)}(\alpha)}=\sup_{\alpha\in[0,\frac{\pi}{2n}]}\frac{\cos\alpha}{\cos\theta_{j_{m}}}.$
For the dual polygonal code, since $k=2$, the Singleton bound implies that there exists a codeword with $c_{(n-1)}=0$, so $h_{n-1}(\mathcal{C})=\infty$. Moreover, the direction ${\bm{u}}(\alpha)$ can be orthogonal to at most one generator vector ${\bm{g}}_{j}$, that is, there is at most one $j$ such that $\cos(\theta_{j}-\alpha)=0$. Hence at most the smallest order statistic $c_{(n-1)}(\alpha)$ can be zero, implying that $\mathcal{C}$ is MDS. In particular, for every $m\leq n-2$, the denominator $c_{(m)}(\alpha)$ is strictly positive for all $\alpha\in[0,\tfrac{\pi}{2n}]$, and the corresponding $m$-height is finite. We therefore restrict attention to $m\leq n-2$.
Theorem III.2.
For the polygonal codes, and $0<m\leq n-2$, the $m$-height achieves its maximum at
$\arg\max_{\alpha\in[0,\frac{\pi}{2n}]}h_{m}(\alpha)=\begin{cases}\alpha=\tfrac{\pi}{2n},&\text{$m$ even},\\[2.0pt] \alpha=0,&\text{$m$ odd}.\end{cases}$
Moreover, the $m$-height is
$h_{m}(\mathcal{C})=\begin{cases}\displaystyle\dfrac{\cos\tfrac{\pi}{2n}}{\cos\bigl((m+1)\tfrac{\pi}{2n}\bigr)},&\text{$m$ even},\\[10.0pt] \displaystyle\dfrac{1}{\cos\bigl((m+1)\tfrac{\pi}{2n}\bigr)},&\text{$m$ odd}.\end{cases}$
Proof.
From Lemma III.1, the ordered term for $0<m\leq n-2$ satisfies
$c_{(m)}(\varphi)=\cos\!\big(A_{(m)}+s_{(m)}\varphi\big),$
where
$A_{(m)}=\frac{\pi}{n}\Bigl\lceil\frac{m}{2}\Bigr\rceil,\qquad s_{(m)}=(-1)^{m},$
and therefore
$h_{m}(\varphi)=\frac{\cos\varphi}{\cos\!\big(A_{(m)}+s_{(m)}\varphi\big)}.$
Taking a derivative and simplifying,
$h_{m}^{\prime}(\varphi)=\frac{s_{(m)}\sin A_{(m)}}{\cos^{2}\!\big(A_{(m)}+s_{(m)}\varphi\big)}.$
Since $A_{(m)}\in(0,\tfrac{\pi}{2}]$, the numerator has the same sign as $s_{(m)}$, and the denominator is strictly positive. Thus,
$\operatorname{sgn}\,h_{m}^{\prime}(\varphi)=\operatorname{sgn}(s_{(m)})=\begin{cases}+1,&m~\text{even},\\ -1,&m~\text{odd}.\end{cases}$
Hence $h_{m}(\varphi)$ is strictly increasing when $m$ is even and strictly decreasing when $m$ is odd, proving
$\arg\max h_{m}=\begin{cases}\varphi=\tfrac{\pi}{2n},&m~\text{even},\\ \varphi=0,&m~\text{odd}.\end{cases}$
Finally, evaluating at the maximizing point:
$h_{m}(\mathcal{C})=\begin{cases}\displaystyle\dfrac{\cos\tfrac{\pi}{2n}}{\cos\bigl((m+1)\tfrac{\pi}{2n}\bigr)},&\text{$m$ even},\\[10.0pt] \displaystyle\dfrac{1}{\cos\bigl((m+1)\tfrac{\pi}{2n}\bigr)},&\text{$m$ odd}.\end{cases}$
This completes the proof. ∎
IV Dual Polyhedral Codes
In this section, we introduce two codes derived from three-dimensional geometric structures: the icosahedron and the dodecahedron. These codes are related to the dual icosahedral code ${\mathcal{C}}_{\text{ico}}^{\perp}$ and dual duodecahedral code ${\mathcal{C}}_{\text{dod}}^{\perp}$ introduced in Examples 5 and 6 of [11], respectively. The codes we analyze differ from those constructions by at most permutations and sign changes of columns and/or rows of the generator matrix, which will not affect the $m$-height profile.
IV-A Dual icosahedral codes
The icosahedron has $12$ vertices, $20$ faces, and $30$ edges. By placing its vertices on the sphere, we obtain $6$ symmetric axes $\{\bm{g}_{\bm{1}},\ldots,\bm{g}_{\bm{6}}\}$, where each axis corresponds to a pair of antipodal vertices. These six axes can be represented by the following matrix:
$G=\begin{bmatrix}0&0&1&1&\varphi&\varphi\\ 1&1&\varphi&-\varphi&0&0\\ \varphi&-\varphi&0&0&1&-1\end{bmatrix},\qquad\varphi=\frac{1+\sqrt{5}}{2}.$
When searching for the optimal information vector $\bm{u}$} that attains the $m$-height, only its direction matters (i.e., its magnitude does not affect the $m$-height.) Therefore, it is sufficient to restrict the analysis to the faces of the solid, which cover all possible directions. Among the $20$ faces, all are symmetric, and furthermore, each face contains subregions that are themselves symmetric. Thus, analyzing a single representative subregion is equivalent to analyzing the entire space.
Motivated by this symmetry, we restrict our attention to the subregion whose direction is closest to ${\bm{g}}_{1}$, as illustrated in Fig. 2. We define the smaller triangle as
$T:=\mathrm{conv}\{{\bm{v}}_{1},{\bm{v}}_{2},{\bm{v}}_{3}\}$
with
${\bm{v}}_{1}:={\bm{g}}_{1},\qquad{\bm{v}}_{2}:=\frac{{\bm{g}}_{1}+{\bm{g}}_{3}}{2},\qquad{\bm{v}}_{3}:=\frac{{\bm{g}}_{1}+{\bm{g}}_{3}+{\bm{g}}_{5}}{3}.$
${\bm{g}}_{1}$${\bm{g}}_{5}$${\bm{g}}_{3}$${\bm{v}}_{2}$${\bm{v}}_{3}$
Figure 2: Icosahedron with a shaded triangular region indicating the fundamental search space.
Any ${\bm{x}}\in T$ can be written in barycentric form
${\bm{x}}(u,v):=u{\bm{v}}_{1}+v{\bm{v}}_{2}+(1-u-v){\bm{v}}_{3}.$
The parameter domain is the standard triangle
$D:=\{(u,v)\in\mathbb{R}^{2}:u\geq 0,\ v\geq 0,\ u+v\leq 1\}.$
Lemma IV.1.
For every ${\bm{x}}\in T$, the six absolute values $|{\bm{x}}\cdot{\bm{g}}_{i}|$, $i=1,\dots,6$, satisfy the global order
$|{\bm{x}}\cdot{\bm{g}}_{1}|\;\geq\;|{\bm{x}}\cdot{\bm{g}}_{3}|\;\geq\;|{\bm{x}}\cdot{\bm{g}}_{5}|\;\geq\;|{\bm{x}}\cdot{\bm{g}}_{4}|\;\geq\;|{\bm{x}}\cdot{\bm{g}}_{2}|\;\geq\;|{\bm{x}}\cdot{\bm{g}}_{6}|.$
Proof.
A direct computation with the parametrization ${\bm{x}}(u,v)$ yields
${\bm{x}}\cdot{\bm{g}}_{1}$ $=\frac{4u+v}{3}+\frac{\sqrt{5}}{2}+\frac{7}{6},$
${\bm{x}}\cdot{\bm{g}}_{3}$ $=\frac{-2u+v}{3}+\frac{\sqrt{5}}{2}+\frac{7}{6},$
${\bm{x}}\cdot{\bm{g}}_{5}$ $=\frac{-2u-2v}{3}+\frac{\sqrt{5}}{2}+\frac{7}{6},$
and
${\bm{x}}\cdot{\bm{g}}_{2}$ $=-\frac{\varphi}{3}\,(2u-v+1),$
${\bm{x}}\cdot{\bm{g}}_{4}$ $=-\frac{\varphi}{3}\,(2u+2v+1),$
${\bm{x}}\cdot{\bm{g}}_{6}$ $=-\frac{\varphi}{3}\,(4u+v-1).$
From these expressions we first obtain, for all $(u,v)\in T$,
${\bm{x}}\cdot{\bm{g}}_{1}>0,\quad{\bm{x}}\cdot{\bm{g}}_{3}>0,\quad{\bm{x}}\cdot{\bm{g}}_{5}>0,\quad{\bm{x}}\cdot{\bm{g}}_{4}<0,\quad{\bm{x}}\cdot{\bm{g}}_{2}\leq 0,$
and ${\bm{x}}\cdot{\bm{g}}_{6}$ changes sign across the line $4u+v=1$.
For ${\bm{g}}_{1},{\bm{g}}_{3},{\bm{g}}_{5}$, we have
${\bm{x}}\cdot{\bm{g}}_{1}-{\bm{x}}\cdot{\bm{g}}_{3}=2u\geq 0,\qquad{\bm{x}}\cdot{\bm{g}}_{3}-{\bm{x}}\cdot{\bm{g}}_{5}=v\geq 0,$
with equality only on the edges $u=0$ and $v=0$, respectively. Since these three quantities are always positive, this implies
$|{\bm{x}}\cdot{\bm{g}}_{1}|\geq|{\bm{x}}\cdot{\bm{g}}_{3}|\geq|{\bm{x}}\cdot{\bm{g}}_{5}|$
for all ${\bm{x}}\in T$.
To compare $|{\bm{x}}\cdot{\bm{g}}_{5}|$ with $|{\bm{x}}\cdot{\bm{g}}_{4}|$, note that ${\bm{x}}\cdot{\bm{g}}_{5}>0$ and ${\bm{x}}\cdot{\bm{g}}_{4}<0$, so
$|{\bm{x}}\cdot{\bm{g}}_{5}|\geq|{\bm{x}}\cdot{\bm{g}}_{4}|\quad\Longleftrightarrow\quad{\bm{x}}\cdot{\bm{g}}_{5}+{\bm{x}}\cdot{\bm{g}}_{4}\geq 0.$
Using the formulas above, one finds
${\bm{x}}\cdot{\bm{g}}_{5}+{\bm{x}}\cdot{\bm{g}}_{4}=\frac{\sqrt{5}+3}{3}\,(1-u-v),$
which is nonnegative on $T$ and becomes zero exactly when $u+v=1$. Similarly, since ${\bm{x}}\cdot{\bm{g}}_{2}\leq 0$ and ${\bm{x}}\cdot{\bm{g}}_{4}<0$, we compare
$|{\bm{x}}\cdot{\bm{g}}_{4}|\geq|{\bm{x}}\cdot{\bm{g}}_{2}|\quad\Longleftrightarrow\quad{\bm{x}}\cdot{\bm{g}}_{4}-{\bm{x}}\cdot{\bm{g}}_{2}\leq 0,$
and a direct computation shows
${\bm{x}}\cdot{\bm{g}}_{4}-{\bm{x}}\cdot{\bm{g}}_{2}=-\varphi\,v,$
which is nonpositive on $T$, with equality only when $v=0$. For $|{\bm{x}}\cdot{\bm{g}}_{2}|$ and $|{\bm{x}}\cdot{\bm{g}}_{6}|$,
$|{\bm{x}}\cdot{\bm{g}}_{2}|\geq|{\bm{x}}\cdot{\bm{g}}_{6}|\ \Longleftrightarrow\ 2u-v+1\ \geq\ |\,4u+v-1\,|.$
Split into two cases:
(i) If $4u+v\geq 1$, then |4u+v1|=4u+v1|4u+v-1|=4u+v-1 and
$2u-v+1\geq 4u+v-1\ \Longleftrightarrow\ 2\geq 2u+2v\ \Longleftrightarrow\ 1\geq u+v,$
which holds on $T$, with equality exactly on the edge $u+v=1$.
(ii) If $4u+v\leq 1$, then |4u+v1|=14uv|4u+v-1|=1-4u-v and
$2u-v+1\geq 1-4u-v\ \Longleftrightarrow\ 6u\geq 0,$
always true, with equality exactly on the edge $u=0$.
Combining (i)–(ii), we obtain $|{\bm{x}}\cdot{\bm{g}}_{2}|\geq|{\bm{x}}\cdot{\bm{g}}_{6}|$ for all $(u,v)\in T$, and equality holds only on $u=0$ or $u+v=1$.
Combining these comparisons yields the global chain
$|{\bm{x}}\cdot{\bm{g}}_{1}|\;\geq\;|{\bm{x}}\cdot{\bm{g}}_{3}|\;\geq\;|{\bm{x}}\cdot{\bm{g}}_{5}|\;\geq\;|{\bm{x}}\cdot{\bm{g}}_{4}|\;\geq\;|{\bm{x}}\cdot{\bm{g}}_{2}|\;\geq\;|{\bm{x}}\cdot{\bm{g}}_{6}|$
for all ${\bm{x}}\in T$, with equalities only in the boundary cases listed above. ∎
Since the code is MDS [11], we have $h_{m}(\mathcal{C})=\infty$ for $m=4,5$. For $m=1,2,3$, the $m$-height optimization reduces to maximizing
$f_{m}(x):=\frac{{\bm{x}}\cdot{\bm{g}}_{1}}{d_{m}({\bm{x}})},\qquad{\bm{x}}\in T,$
where the denominator is given by
$d_{1}({\bm{x}})={\bm{x}}\cdot{\bm{g}}_{3},\qquad d_{2}({\bm{x}})={\bm{x}}\cdot{\bm{g}}_{5},\qquad d_{3}({\bm{x}})=-\,{\bm{x}}\cdot{\bm{g}}_{4}.$
Theorem IV.2.
Let ${\bm{x}}\in T$ and let $f_{1},f_{2},f_{3}$ be defined as above. Then the maximizers over $T$ and the corresponding maximal values are:
  1. 1.
    $\arg\max_{{\bm{x}}\in T}f_{1}({\bm{x}})={\bm{v}}_{1}$, and $f_{1}({\bm{g}}_{1})=\sqrt{5}$.
  2. 2.
    $\arg\max_{{\bm{x}}\in T}f_{2}({\bm{x}})={\bm{v}}_{1}$, and $f_{2}({\bm{g}}_{1})=\sqrt{5}$.
  3. 3.
    $\arg\max_{{\bm{x}}\in T}f_{3}({\bm{x}})={\bm{v}}_{3}$, and $f_{3}({\bm{v}}_{3})=2+\sqrt{5}$.
Proof.
We prove each claim by analyzing the monotonicity of the corresponding objective on the domain $D$.
(1) For $f_{1}$. For $(u,v)\in D$,
$f_{1}(u,v)=\frac{4u+v+\frac{3\sqrt{5}}{2}+\frac{7}{2}}{-2u-2v+\frac{3\sqrt{5}}{2}+\frac{7}{2}}.$
A direct differentiation simplifies to
$\partial_{u}f_{1}(u,v)$ $=\frac{-6v+9\sqrt{5}+21}{\bigl(-2u-2v+\frac{3\sqrt{5}}{2}+\frac{7}{2}\bigr)^{2}}>0,$
$\partial_{v}f_{1}(u,v)$ $=\frac{6u+\frac{9\sqrt{5}}{2}+\frac{21}{2}}{\bigl(-2u-2v+\frac{3\sqrt{5}}{2}+\frac{7}{2}\bigr)^{2}}>0.$
Hence $f_{1}$ is increasing in both $u$ and $v$ on $D$, so the maximum must lie on the edge $u+v=1$. Restricting to this edge,
$f_{1}(u,1-u)=\frac{3u+1+\frac{3\sqrt{5}}{2}+\frac{7}{2}}{-2+\frac{3\sqrt{5}}{2}+\frac{7}{2}},$
which is strictly increasing in $u$. Therefore the unique maximizer is $u=1$, i.e., $(u,v)=(1,0)$, corresponding to ${\bm{x}}={\bm{v}}_{1}={\bm{g}}_{1}$. Evaluating yields $f_{1}({\bm{g}}_{1})=\sqrt{5}$.
(2) For $f_{2}$. For $(u,v)\in D$,
$f_{2}(u,v)=\frac{8u+2v+3\sqrt{5}+7}{-4u+2v+3\sqrt{5}+7}.$
Differentiation gives $\partial_{u}f_{2}(u,v)>0$ and $\partial_{v}f_{2}(u,v)\leq 0$ on $D$. Thus, for any $(u,v)\in D$ we have $f_{2}(u,0)\geq f_{2}(u,v)$, and then $f_{2}(1,0)\geq f_{2}(u,0)$. Hence the maximum is attained at $(u,v)=(1,0)$, i.e., ${\bm{x}}={\bm{v}}_{1}={\bm{g}}_{1}$, and $f_{2}({\bm{g}}_{1})=\sqrt{5}$.
(3) For $f_{3}$. For $(u,v)\in D$,
$f_{3}(u,v)=\frac{8u+2v+3\sqrt{5}+7}{(1+\sqrt{5})(2u+2v+1)}.$
Since $(1+\sqrt{5})^{-1}>0$, it suffices to maximize
$g(u,v)=\frac{8u+2v+3\sqrt{5}+7}{2u+2v+1}.$
One checks that $\partial_{u}g(u,v)<0$ and $\partial_{v}g(u,v)<0$ on $D$. Hence $g$ (and thus $f_{3}$) is strictly decreasing as either $u$ or $v$ increases, so the maximum over $D$ is attained at $(u,v)=(0,0)$, corresponding to ${\bm{x}}={\bm{v}}_{3}$. Evaluating yields $f_{3}({\bm{v}}_{3})=2+\sqrt{5}$.
This completes the proof. ∎
IV-B Dual dodecahedral codes
We now consider the geometric code derived from the regular dodecahedron. The dodecahedron has $20$ vertices, $12$ faces, and $30$ edges. Placing its vertices on the sphere yields $10$ symmetric axes $\{{\bm{g}}_{1},\ldots,{\bm{g}}_{10}\}$ (each axis connects a pair of antipodal vertices), represented by
$G=\begin{bmatrix}1&1&1&1&0&0&\varphi^{-1}&\varphi^{-1}&\varphi&\varphi\\ 1&1&-1&-1&\varphi&\varphi&0&0&\varphi^{-1}&-\varphi^{-1}\\ 1&-1&1&-1&\varphi^{-1}&-\varphi^{-1}&\varphi&-\varphi&0&0\end{bmatrix}$.
Similar to the icosahedral case, symmetry allows us to restrict attention to a fundamental triangular region on a single dodecahedral face, as illustrated in Fig. 3, with vertices
${\bm{x}}_{A}:={\bm{g}}_{1},\ \ {\bm{x}}_{B}:=\frac{{\bm{g}}_{1}+{\bm{g}}_{5}}{2},\ \ {\bm{x}}_{C}:=\frac{{\bm{g}}_{1}+{\bm{g}}_{2}+{\bm{g}}_{5}+{\bm{g}}_{6}+{\bm{g}}_{9}}{5}.$
Let $T^{\prime}=\mathrm{conv}\{{\bm{x}}_{A},{\bm{x}}_{B},{\bm{x}}_{C}\},{\bm{x}}(u,v)=u\,{\bm{x}}_{A}+v\,{\bm{x}}_{B}+(1-u-v)\,{\bm{x}}_{C},$ where $u,v\geq 0$ and $u+v\leq 1$, and define $\beta_{j}(x):=|{\bm{x}}\cdot{\bm{g}}_{j}|$. Let $\beta_{[j]}(x)$ be the order statistics of $\beta_{j}(x)$.
${\bm{g}}_{1}$${\bm{g}}_{9}$${\bm{g}}_{2}$${\bm{g}}_{6}$${\bm{g}}_{5}$${\bm{x}}_{B}$${\bm{x}}_{C}$
Figure 3: Dodecahedron with a shaded triangular region indicating the fundamental search space.
Lemma IV.3.
For any ${\bm{x}}\in T^{\prime}$ we have the following inequalities:
$\beta_{1}\geq\beta_{5}\geq\beta_{9}\geq\max\{\beta_{6},\beta_{7}\},$ (1)
$\beta_{6}\geq\beta_{2},\qquad\beta_{6}\geq\beta_{4},\qquad\beta_{7}\geq\beta_{4},$ (2)
$\beta_{4}\geq\max\{\beta_{3},\beta_{8},\beta_{10}\},\qquad\beta_{2}\geq\max\{\beta_{3},\beta_{8},\beta_{10}\},$ (3)
$\beta_{3}\leq\beta_{10}.$ (4)
Proof.
We first compute
${\bm{x}}\cdot{\bm{g}}_{1}$ $=\Bigl(2-\frac{2\sqrt{5}}{5}\Bigr)u+\Bigl(\frac{1}{2}+\frac{\sqrt{5}}{10}\Bigr)v+\Bigl(1+\frac{2\sqrt{5}}{5}\Bigr),$
${\bm{x}}\cdot{\bm{g}}_{5}$ $=\Bigl(-1+\frac{3\sqrt{5}}{5}\Bigr)u+\Bigl(\frac{1}{2}+\frac{\sqrt{5}}{10}\Bigr)v+\Bigl(1+\frac{2\sqrt{5}}{5}\Bigr),$
${\bm{x}}\cdot{\bm{g}}_{9}$ $=\Bigl(-1+\frac{3\sqrt{5}}{5}\Bigr)u+\Bigl(-\frac{1}{2}+\frac{\sqrt{5}}{10}\Bigr)v+\Bigl(1+\frac{2\sqrt{5}}{5}\Bigr),$
${\bm{x}}\cdot{\bm{g}}_{2}$ $=\Bigl(-\frac{2\sqrt{5}}{5}\Bigr)u+\Bigl(-\frac{2\sqrt{5}}{5}\Bigr)v+\Bigl(1+\frac{2\sqrt{5}}{5}\Bigr),$
${\bm{x}}\cdot{\bm{g}}_{6}$ $=\Bigl(-\frac{2\sqrt{5}}{5}\Bigr)u+\Bigl(-\frac{2\sqrt{5}}{5}+\frac{\sqrt{5}-1}{2}\Bigr)v+\Bigl(1+\frac{2\sqrt{5}}{5}\Bigr),$
${\bm{x}}\cdot{\bm{g}}_{7}$ $=\Bigl(\frac{4\sqrt{5}}{5}\Bigr)u+\Bigl(\frac{5+3\sqrt{5}}{10}\Bigr)v+\frac{\sqrt{5}}{5},$
${\bm{x}}\cdot{\bm{g}}_{4}$ $=\Bigl(-1+\frac{\sqrt{5}}{5}\Bigr)u-\Bigl(\frac{1}{2}+\frac{3\sqrt{5}}{10}\Bigr)v-\frac{\sqrt{5}}{5},$
${\bm{x}}\cdot{\bm{g}}_{10}$ $=\Bigl(1-\frac{\sqrt{5}}{5}\Bigr)u-\frac{\sqrt{5}}{5}\,v+\frac{\sqrt{5}}{5},$
${\bm{x}}\cdot{\bm{g}}_{3}$ $=\Bigl(1+\frac{\sqrt{5}}{5}\Bigr)u+\frac{\sqrt{5}}{5}\,v-\frac{\sqrt{5}}{5},$
${\bm{x}}\cdot{\bm{g}}_{8}$ $=\frac{\sqrt{5}}{5}-\Bigl(1+\frac{\sqrt{5}}{5}\Bigr)(u+v).$
On $T^{\prime}$, one checks that ${\bm{x}}\cdot{\bm{g}}_{4}\leq 0$ everywhere, while
${\bm{x}}\cdot{\bm{g}}_{j}\geq 0\quad\text{for }j\in\{1,2,5,6,7,9,10\}.$
The only inner products that may change sign on $T^{\prime}$ are ${\bm{x}}\cdot{\bm{g}}_{3}$ and ${\bm{x}}\cdot{\bm{g}}_{8}$. Therefore, we may drop absolute values for ${\bm{x}}\cdot{\bm{g}}_{j}$ with $j\in\{1,2,5,6,7,9,10\}$, keep a minus sign for $|{\bm{x}}\cdot{\bm{g}}_{4}|=-({\bm{x}}\cdot{\bm{g}}_{4})$, and retain absolute values for $|{\bm{x}}\cdot{\bm{g}}_{3}|$ and $|{\bm{x}}\cdot{\bm{g}}_{8}|$.
Proof of (1). Compute the differences:
$({\bm{x}}\cdot{\bm{g}}_{1})-({\bm{x}}\cdot{\bm{g}}_{5})=(3-\sqrt{5})\,u\ \geq\ 0,\ \ \ ({\bm{x}}\cdot{\bm{g}}_{5})-({\bm{x}}\cdot{\bm{g}}_{9})=v\ \geq\ 0,$
$({\bm{x}}\cdot{\bm{g}}_{9})-({\bm{x}}\cdot{\bm{g}}_{6})=\Bigl(\frac{5-\sqrt{5}}{5}\Bigr)u+\frac{\sqrt{5}-1}{2}\,v\ \geq 0,$
$({\bm{x}}\cdot{\bm{g}}_{9})-({\bm{x}}\cdot{\bm{g}}_{7})=\Bigl(1+\frac{\sqrt{5}}{5}\Bigr)\bigl(1-u-v\bigr)\ \geq 0.$
which gives $\beta_{1}\geq\beta_{5}\geq\beta_{9}\geq\max\{\beta_{6},\beta_{7}\}$.
Proof of (2). First,
$({\bm{x}}\cdot{\bm{g}}_{6})-({\bm{x}}\cdot{\bm{g}}_{2})=\frac{\sqrt{5}-1}{2}\,v\ \geq\ 0\quad\Longrightarrow\quad\beta_{6}\geq\beta_{2}.$
Next,
$\beta_{6}-\beta_{4}=({\bm{x}}\cdot{\bm{g}}_{6})+({\bm{x}}\cdot{\bm{g}}_{4})=1+\frac{\sqrt{5}}{5}-\Bigl(1+\frac{\sqrt{5}}{5}\Bigr)(u+v)\ \geq\ 0,$
so $\beta_{6}\geq\beta_{4}$. Also,
$\beta_{7}-\beta_{4}=({\bm{x}}\cdot{\bm{g}}_{7})+({\bm{x}}\cdot{\bm{g}}_{4})=(\sqrt{5}-1)\,u\ \geq\ 0.$
so $\beta_{7}\geq\beta_{4}$.
Proof of (3). We show $|{\bm{x}}\cdot{\bm{g}}_{j}|\leq\beta_{4}$ and $|{\bm{x}}\cdot{\bm{g}}_{j}|\leq\beta_{2}$ for $j\in\{3,8,10\}$. For $j=10$ (no absolute needed):
$\beta_{4}-\beta_{10}=-({\bm{x}}\cdot{\bm{g}}_{4})-({\bm{x}}\cdot{\bm{g}}_{10})=\frac{1+\sqrt{5}}{2}\,v\ \geq 0,$
$\beta_{2}-\beta_{10}=1+\frac{\sqrt{5}}{5}-\Bigl(1+\frac{\sqrt{5}}{5}\Bigr)u-\frac{\sqrt{5}}{5}\,v\ \geq 0.$
For $j=3$, it suffices to check $\beta_{4}\pm({\bm{x}}\cdot{\bm{g}}_{3})\geq 0$ and $\beta_{2}\pm({\bm{x}}\cdot{\bm{g}}_{3})\geq 0$:
$\beta_{4}-({\bm{x}}\cdot{\bm{g}}_{3})=\frac{2\sqrt{5}}{5}(1-u)+\frac{5+\sqrt{5}}{10}v\ \geq 0,$
$\beta_{4}+({\bm{x}}\cdot{\bm{g}}_{3})=2u+\Bigl(\frac{1}{2}+\frac{\sqrt{5}}{2}\Bigr)v\ \geq 0,$
$\beta_{2}-({\bm{x}}\cdot{\bm{g}}_{3})=1+\frac{3\sqrt{5}}{5}-\Bigl(1+\frac{3\sqrt{5}}{5}\Bigr)u-\frac{3\sqrt{5}}{5}v\ \geq 0,$
$\beta_{2}+({\bm{x}}\cdot{\bm{g}}_{3})=1+\frac{\sqrt{5}}{5}+\Bigl(1-\frac{\sqrt{5}}{5}\Bigr)u-\frac{\sqrt{5}}{5}v\ \geq 0.$
For $j=8$, similarly check $\beta_{4}\pm({\bm{x}}\cdot{\bm{g}}_{8})\geq 0$ and $\beta_{2}\pm({\bm{x}}\cdot{\bm{g}}_{8})\geq 0$:
$\beta_{4}-({\bm{x}}\cdot{\bm{g}}_{8})=2u+\Bigl(\frac{1}{2}+\frac{3\sqrt{5}}{10}\Bigr)v+\frac{\sqrt{5}}{5}\ \geq 0,$
$\beta_{4}+({\bm{x}}\cdot{\bm{g}}_{8})=\frac{2\sqrt{5}}{5}-\frac{2\sqrt{5}}{5}u-\frac{5-\sqrt{5}}{10}\,v\ \geq 0,$
$\beta_{2}-({\bm{x}}\cdot{\bm{g}}_{8})=1+\frac{3\sqrt{5}}{5}-\Bigl(1-\frac{\sqrt{5}}{5}\Bigr)u-\Bigl(1-\frac{3\sqrt{5}}{5}\Bigr)v\ \geq 0,$
$\beta_{2}+({\bm{x}}\cdot{\bm{g}}_{8})=1+\frac{\sqrt{5}}{5}+\Bigl(1+\frac{\sqrt{5}}{5}\Bigr)u+\frac{3\sqrt{5}}{5}v\ \geq 0.$
Thus $\beta_{4}\geq\max\{\beta_{3},\beta_{8},\beta_{10}\}$ and $\beta_{2}\geq\max\{\beta_{3},\beta_{8},\beta_{10}\}$.
Proof of (4). Since $\beta_{10}={\bm{x}}\cdot{\bm{g}}_{10}\geq 0$, it suffices to show $\beta_{10}\pm({\bm{x}}\cdot{\bm{g}}_{3})\geq 0$:
$\beta_{10}-({\bm{x}}\cdot{\bm{g}}_{3})$ $=\frac{2\sqrt{5}}{5}(1-u-v)\ \geq 0,$
$\beta_{10}+({\bm{x}}\cdot{\bm{g}}_{3})$ $=2u\ \geq 0.$
hence $|{\bm{x}}\cdot{\bm{g}}_{3}|\leq{\bm{x}}\cdot{\bm{g}}_{10}$, i.e., $\beta_{3}\leq\beta_{10}$. ∎
Corollary IV.4.
Let ${\bm{x}}\in T^{\prime}$ and let $\beta_{[k]}({\bm{x}})$ be the order statistics of $\beta_{j}({\bm{x}})$. Under the inequalities (1)–(4), the index attaining rank $k$ can only belong to:
$k=1:\{1\},\quad k=2:\{5\},\quad k=3:\{9\},\quad k=4:\{6,7\},\quad$
$k=5:\{2,6,7\},\quad k=6:\{2,4,7\},\quad k=7:\{2,4\},$
$k=8:\{8,10\},\quad k=9:\{3,8,10\},\quad k=10:\{3,8\}.$
Proof.
From (1)–(4), we know that ranks $1,2,3$ are attained uniquely by $\{1\},\{5\},\{9\}$, and the remaining ranks are determined by $\{\beta_{2},\beta_{3},\beta_{4},\beta_{6},\beta_{7},\beta_{8},\beta_{10}\}$.
From (3), we have
$\max\{\beta_{3},\beta_{8},\beta_{10}\}\leq\min\{\beta_{2},\beta_{4}\},$
so $\{3,8,10\}$ must occupy the bottom three ranks $k=8,9,10$, and $\{2,4,6,7\}$ must occupy ranks $k=4,5,6,7$.
Within $\{2,4,6,7\}$, inequality (2) gives
$\beta_{6}\geq\beta_{2},\qquad\beta_{6}\geq\beta_{4},\qquad\beta_{7}\geq\beta_{4}.$
Hence $\beta_{2}$ cannot be the largest among $\{\beta_{2},\beta_{4},\beta_{6},\beta_{7}\}$ since $\beta_{6}\geq\beta_{2}$, and $\beta_{4}$ cannot be the largest since $\beta_{6}\geq\beta_{4}$; therefore the largest must be attained at index $6$ or $7$, i.e., $k=4:\{6,7\}$. Moreover, since both $\beta_{6}$ and $\beta_{7}$ dominate $\beta_{4}$, index $4$ cannot be the second-largest, so the second-largest must lie in $\{2,6,7\}$, i.e., $k=5:\{2,6,7\}$. Next, because $\beta_{6}\geq\beta_{2}$ and $\beta_{6}\geq\beta_{4}$, index $6$ cannot be the third- or fourth-largest within this subset; thus the third-largest must lie in $\{2,4,7\}$, i.e., $k=6:\{2,4,7\}$. Finally, since $\beta_{7}\geq\beta_{4}$, index $7$ cannot be the smallest among the four, so the smallest must lie in $\{2,4\}$, i.e., $k=7:\{2,4\}$.
Within the bottom group $\{3,8,10\}$, inequality (4) gives $\beta_{3}\leq\beta_{10}$, so index $3$ cannot be the largest among $\{3,8,10\}$ and index $10$ cannot be the smallest among $\{3,8,10\}$. Hence
$k=8:\{8,10\},\qquad k=9:\{3,8,10\},\qquad k=10:\{3,8\}.$
Lemma IV.5.
For each $j\in\{2,4,5,6,7,9,10\}$, define
$f_{j}(u,v):=\frac{{\bm{x}}(u,v)\cdot{\bm{g}}_{1}}{|{\bm{x}}(u,v)\cdot{\bm{g}}_{j}|},\qquad(u,v)\in T^{\prime}.$
Then $f_{j}$ has no stationary point in $T^{\prime}$. In fact, on $T^{\prime}$ we have
$\partial_{u}f_{2}>0,\ \partial_{v}f_{2}>0;\qquad\partial_{u}f_{5}>0,\ \partial_{v}f_{5}\leq 0;$
$\partial_{u}f_{6}>0,\ \partial_{v}f_{6}>0;\qquad\partial_{u}f_{7}<0,\ \partial_{v}f_{7}<0;$
$\partial_{u}f_{9}>0,\ \partial_{v}f_{9}>0;\qquad\partial_{u}f_{10}<0,\ \partial_{v}f_{10}>0;$
and for $j=4$, $\partial_{v}f_{4}<0.$
Proof.
On $T^{\prime}$, we have ${\bm{x}}\cdot{\bm{g}}_{4}<0$ and ${\bm{x}}\cdot{\bm{g}}_{j}>0$ for $j\in\{2,5,6,7,9,10\}$. Hence $|{\bm{x}}\cdot{\bm{g}}_{4}|=-({\bm{x}}\cdot{\bm{g}}_{4})$ and $|{\bm{x}}\cdot{\bm{g}}_{j}|={\bm{x}}\cdot{\bm{g}}_{j}$ for $j\in\{2,5,6,7,9,10\}$.
Write $N:={\bm{x}}\cdot{\bm{g}}_{1}=a_{1}u+b_{1}v+c_{1}$ and $D_{j}:=|{\bm{x}}\cdot{\bm{g}}_{j}|=a_{j}u+b_{j}v+c_{j}$. We have
$\partial_{u}f_{j}=\frac{a_{1}D_{j}-a_{j}N}{D_{j}^{2}},\qquad\partial_{v}f_{j}=\frac{b_{1}D_{j}-b_{j}N}{D_{j}^{2}}.$
Substituting the explicit formulas of ${\bm{x}}\cdot{\bm{g}}_{1}$ and ${\bm{x}}\cdot{\bm{g}}_{j}$ yields the following:
$\partial_{u}f_{2}$ $=\frac{(10+4\sqrt{5})+(5-3\sqrt{5})v}{5\,D_{2}^{2}}>0,$
$\partial_{v}f_{2}$ $=\frac{(15+7\sqrt{5})+(-10+6\sqrt{5})u}{10\,D_{2}^{2}}>0;$
$\partial_{u}f_{5}$ $=\frac{(5+\sqrt{5})+(5-\sqrt{5})v}{5\,D_{5}^{2}}>0,$
$\partial_{v}f_{5}$ $=\frac{(\sqrt{5}-5)u}{5\,D_{5}^{2}}<0;$
$\partial_{u}f_{6}$ $=\frac{(10+4\sqrt{5})+(-5+3\sqrt{5})v}{5\,D_{6}^{2}}>0,$
$\partial_{v}f_{6}$ $=\frac{(5+2\sqrt{5})+(5-3\sqrt{5})u}{5\,D_{6}^{2}}>0;$
$\partial_{u}f_{7}$ $=-\frac{2(5+\sqrt{5})}{5\,D_{7}^{2}}<0,$
$\partial_{v}f_{7}$ $=-\frac{5+2\sqrt{5}}{5\,D_{7}^{2}}<0;$
$\partial_{u}f_{9}$ $=\frac{(5+\sqrt{5})+(-5+\sqrt{5})v}{5\,D_{9}^{2}}>0,$
$\partial_{v}f_{9}$ $=\frac{(5+2\sqrt{5})+(5-\sqrt{5})u}{5\,D_{9}^{2}}>0;$
$\partial_{u}f_{10}$ $=-\frac{(5-\sqrt{5})+2\sqrt{5}\,v}{5\,D_{10}^{2}}<0,$
$\partial_{v}f_{10}$ $=\frac{(5+3\sqrt{5})+4\sqrt{5}\,u}{10\,D_{10}^{2}}>0.$
For $j=4$ we use $D_{4}=|{\bm{x}}\cdot{\bm{g}}_{4}|=-({\bm{x}}\cdot{\bm{g}}_{4})=\bigl(1-\frac{\sqrt{5}}{5}\bigr)u+\bigl(\frac{1}{2}+\frac{3\sqrt{5}}{10}\bigr)v+\frac{\sqrt{5}}{5}$, and obtain
$\partial_{v}f_{4}=-\frac{2\sqrt{5}\,u+(5+2\sqrt{5})}{5\,D_{4}^{2}}<0.$
In each case, $\nabla f_{j}$ cannot be zero on $T^{\prime}$, so $f_{j}$ has no stationary point in $T^{\prime}$. ∎
Corollary IV.6.
Let ${\bm{x}}\in T^{\prime}$. For $m=1,2$, the corresponding $m$-height objective functions are, respectively,
$f_{5}({\bm{x}}):=\frac{{\bm{x}}\cdot{\bm{g}}_{1}}{{\bm{x}}\cdot{\bm{g}}_{5}}\qquad and\qquad f_{9}({\bm{x}}):=\frac{{\bm{x}}\cdot{\bm{g}}_{1}}{{\bm{x}}\cdot{\bm{g}}_{9}}.$
Moreover:
  1. 1)
    The maximizer of $f_{5}$ over $T^{\prime}$ is ${\bm{x}}={\bm{x}}_{A}={\bm{g}}_{1}$, and
    $f_{5}({\bm{g}}_{1})=\frac{3}{\sqrt{5}}.$
  2. 2)
    The maximizer of $f_{9}$ over $T^{\prime}$ is ${\bm{x}}={\bm{x}}_{B}=\tfrac{{\bm{g}}_{1}+{\bm{g}}_{5}}{2}$, and
    $f_{9}({\bm{x}}_{B})=\varphi.$
Proof.
(1) Case $m=1$. From the proof of Lemma IV.5, we have
$\partial_{u}f_{5}(u,v)\;>\;0,\qquad\partial_{v}f_{5}(u,v)\;<\;0.$
Hence $f_{5}$ is increasing in $u$ and decreasing in $v$ on $T^{\prime}$, so the maximizer is attained at $(u,v)=(1,0)$, i.e., ${\bm{x}}={\bm{x}}_{A}={\bm{g}}_{1}$. Evaluating gives
$f_{5}({\bm{g}}_{1})=\frac{{\bm{g}}_{1}\cdot{\bm{g}}_{1}}{{\bm{g}}_{1}\cdot{\bm{g}}_{5}}=\frac{3}{\sqrt{5}}.$
(2) Case $m=2$. We have
$\partial_{u}f_{9}(u,v)\;>\;0,\qquad\partial_{v}f_{9}(u,v)\;>\;0.$
Thus $f_{9}$ is increasing in both $u$ and $v$, and by monotonicity the maximizer lies on the edge $u+v=1$. Along this edge,
$f_{9}(u,1-u)=\frac{(3-\sqrt{5})u+(3+\sqrt{5})}{(\sqrt{5}-1)u+(1+\sqrt{5})}$
is strictly decreasing in $u$, so the unique maximizer is $(u,v)=(0,1)$, i.e., ${\bm{x}}={\bm{x}}_{B}=\tfrac{{\bm{g}}_{1}+{\bm{g}}_{5}}{2}$. Evaluating gives
$f_{9}({\bm{x}}_{B})=\frac{{\bm{x}}_{B}\cdot{\bm{g}}_{1}}{{\bm{x}}_{B}\cdot{\bm{g}}_{9}}=\varphi.$
This completes the proof. ∎
Theorem IV.7.
For each $m\in\{3,4,5,6,7\}$, any maximizer of the corresponding $m$-height over $T^{\prime}$ must lie in the candidate set
$\mathcal{S}=\bigl\{(u,v)\in T^{\prime}:\ (1,0),\ (0,0),\ (0,1),\bigr$.
$\bigl.\bigl(0,\tfrac{1+3\sqrt{5}}{11}\bigr),\ \bigl(\tfrac{\varphi}{3},0\bigr),\ \bigl(0,\,2\sqrt{5}-4\bigr)\bigr\}.$
Proof.
We treat the three cases $m\in\{3,4\}$, $m\in\{5,6\}$, and $m=7$ separately.
Case I: $m=3,4$ (the denominator is in $\{{\bm{g}}_{2},{\bm{g}}_{6},{\bm{g}}_{7}\}$). By Corollary IV.4, for $m=3,4$, the (m+1)(m\!+\!1)-th largest projection must be attained at index $2$, 66, or $7$.
Define
$L_{i,j}:=\{{\bm{x}}\in T^{\prime}:\ |{\bm{x}}\cdot{\bm{g}}_{i}|=|{\bm{x}}\cdot{\bm{g}}_{j}|\}.$
Since ${\bm{x}}\cdot{\bm{g}}_{2},\ {\bm{x}}\cdot{\bm{g}}_{6},\ {\bm{x}}\cdot{\bm{g}}_{7}\geq 0$ on $T^{\prime}$, we drop the absolute values and write $L_{i,j}=\{{\bm{x}}\in T^{\prime}:{\bm{x}}\cdot{\bm{g}}_{i}={\bm{x}}\cdot{\bm{g}}_{j}\}$ for $i,j\in\{2,6,7\}$. The switching boundaries among these three denominators are given by (we do not need to consider the switching boundary $L_{2,6}$ since by Lemma IV.3, ${\bm{x}}\cdot{\bm{g}}_{6}\geq{\bm{x}}\cdot{\bm{g}}_{2}$)
$L_{2,7}:\ {\bm{x}}\cdot{\bm{g}}_{2}={\bm{x}}\cdot{\bm{g}}_{7}\iff 12\sqrt{5}\,u+(7\sqrt{5}+5)\,v=10+2\sqrt{5},$
$L_{6,7}:\ {\bm{x}}\cdot{\bm{g}}_{6}={\bm{x}}\cdot{\bm{g}}_{7}\iff 6\sqrt{5}\,u+(5+\sqrt{5})\,v=5+\sqrt{5}.$
By Lemma IV.5, none of $f_{2}$, $f_{6}$, or $f_{7}$ has a stationary point in $T^{\prime}$. The switching boundaries partition $T^{\prime}$ into three subregions, as illustrated in Fig. 4. Within each subregion, the ordering among the denominators is fixed, and the relevant denominator is ${\bm{g}}_{j}$ for some $j\in\{2,6,7\}$. Consequently, the $m$-height optimization reduces to maximizing a single ratio $f_{j}(u,v)$ over that subregion. We thus analyze the three subregions separately, maximizing $f_{2}(u,v)$, $f_{6}(u,v)$, or $f_{7}(u,v)$ on the corresponding region.
[…]
Figure 4: Triangle $T^{\prime}$ in $(u,v)$-coordinates and switching lines $L_{2,7}$ and $L_{6,7}$, partitioning $T^{\prime}$ into three subregions.
(A) Maximizers in $T_{1}$. From Lemma IV.5, both $\partial_{u}f_{2}$ and $\partial_{v}f_{2}$ are strictly positive. Hence $f_{2}$ is strictly increasing in both $u$ and $v$, and its maximum over $T_{1}$ must be attained on the edge $L_{2,7}$. Evaluating $f_{2}$ along this edge shows that
$f_{2}\Bigl(u,\frac{10+2\sqrt{5}-12\sqrt{5}\,u}{7\sqrt{5}+5}\Bigr)=\frac{(13-5\sqrt{5})u+(13+6\sqrt{5})}{(4\sqrt{5}-6)u+(5+4\sqrt{5})},$
with derivative
$\frac{d}{du}(\cdot)=\frac{11(\sqrt{5}-7)}{\bigl((4\sqrt{5}-6)u+(5+4\sqrt{5})\bigr)^{2}}<0.$
Hence the maximum is attained at $u=0$, i.e.,
$(u,v)=\Bigl(0,\ \frac{1+3\sqrt{5}}{11}\Bigr).$
For $f_{6}$, both $\partial_{u}f_{6}$ and $\partial_{v}f_{6}$ are strictly positive. Evaluating $f_{6}$ along $L_{2,7}$ shows that
$f_{6}\bigl(u,v(u)\bigr)=\frac{(6\sqrt{5}-10)u+(25+11\sqrt{5})}{(4\sqrt{5}-20)u+(15+9\sqrt{5})}.$
Differentiating yields
$\frac{d}{du}f_{6}\bigl(u,v(u)\bigr)=\frac{40(10+3\sqrt{5})}{\bigl((4\sqrt{5}-20)u+(15+9\sqrt{5})\bigr)^{2}}>0.$
Hence $f_{6}$ is strictly increasing along $L_{2,7}$ inside $T^{\prime}$, so its maximum on this segment is attained at the endpoint with largest $u$, namely where $v=0$:
$(u,v)=\Bigl(\frac{10+2\sqrt{5}}{12\sqrt{5}},\,0\Bigr)=\Bigl(\frac{1+\sqrt{5}}{6},\,0\Bigr).$
For $f_{7}$, both partial derivatives satisfy $\partial_{u}f_{7}<0$ and $\partial_{v}f_{7}<0$ on $T^{\prime}$, so $f_{7}$ is strictly decreasing in both variables. Hence its maximum over $T_{1}$ is attained at the vertex $(u,v)=(0,0)$.
(B) Maximizers in $T_{2}$.
For $f_{2}$, both $\partial_{u}f_{2}$ and $\partial_{v}f_{2}$ are strictly positive. Evaluating $f_{2}$ along $L_{6,7}$ shows that
$f_{2}\Bigl(u,\ 1+\tfrac{3}{2}(1-\sqrt{5})u\Bigr)=\frac{(4-2\sqrt{5})u+(3+\sqrt{5})}{2\bigl((3-\sqrt{5})u+1\bigr)}.$
Differentiating gives
$\frac{d}{du}(\cdot)=\frac{-4\sqrt{5}}{\bigl(2((3-\sqrt{5})u+1)\bigr)^{2}}<0,$
hence the maximum is attained at $u=0$, i.e., at $(u,v)=(0,1)$.
For $f_{6}$, both $\partial_{u}f_{6}$ and $\partial_{v}f_{6}$ are strictly positive. Evaluating $f_{6}$ along $L_{6,7}$ shows that
$f_{6}\Bigl(u,\ 1+\tfrac{3}{2}(1-\sqrt{5})u\Bigr)=\frac{(4-2\sqrt{5})u+(3+\sqrt{5})}{(\sqrt{5}-3)u+(1+\sqrt{5})}.$
with derivative
$\frac{d}{du}(\cdot)=\frac{10(1+\sqrt{5})}{\bigl((\sqrt{5}-3)u+(1+\sqrt{5})\bigr)^{2}}>0.$
Thus the maximum on $L_{6,7}$ is attained at the largest feasible $u$, which occurs at $v=0$, i.e.,
$(u,v)=\Bigl(\frac{1+\sqrt{5}}{6},\ 0\Bigr).$
For $f_{7}$, both $\partial_{u}f_{7}$ and $\partial_{v}f_{7}$ are strictly negative. Evaluating $f_{7}$ along $L_{2,7}$ shows that
$f_{7}\Bigl(u,\frac{10+2\sqrt{5}-12\sqrt{5}\,u}{7\sqrt{5}+5}\Bigr)=\frac{(6\sqrt{5}-10)u+(25+11\sqrt{5})}{(10-2\sqrt{5})u+(15+5\sqrt{5})}.$
Its derivative is
$\frac{d}{du}(\cdot)=-\frac{20(7+\sqrt{5})}{\bigl((10-2\sqrt{5})u+(15+5\sqrt{5})\bigr)^{2}}<0.$
Hence the maximum on this segment is attained at the endpoint with smallest $u$, i.e.,
$(u,v)=\Bigl(0,\ \frac{1+3\sqrt{5}}{11}\Bigr).$
(C) Maximizers in $T_{3}$.
For $f_{2}$, both $\partial_{u}f_{2}$ and $\partial_{v}f_{2}$ are strictly positive. Evaluating $f_{2}$ along edge $AB$ shows that
$f_{2}(u,1-u)=\bigl(\frac{3}{2}-\frac{\sqrt{5}}{2}\bigr)u+\bigl(\frac{3}{2}+\frac{\sqrt{5}}{2}\bigr).$
Differentiating yields
$\frac{d}{du}(\cdot)=\frac{3}{2}-\frac{\sqrt{5}}{2}>0,$
so $f_{2}$ is strictly increasing along $AB$. Hence the maximum on $AB$ is attained at $u=1$, i.e., at the vertex $A=(1,0)$.
For $f_{6}$, both $\partial_{u}f_{6}$ and $\partial_{v}f_{6}$ are strictly positive. Evaluating $f_{6}$ along edge $AB$ shows that
$f_{6}(u,1-u)=\frac{(3-\sqrt{5})u+(3+\sqrt{5})}{(1-\sqrt{5})u+(1+\sqrt{5})}.$
Its derivative is
$\frac{d}{du}f_{6}(u,1-u)=\frac{4\sqrt{5}}{\bigl((1-\sqrt{5})u+(1+\sqrt{5})\bigr)^{2}}>0,$
so $f_{6}(u,1-u)$ is strictly increasing on $u$. Therefore the maximum on edge $AB$ is attained at $u=1$ (i.e., $(u,v)=(1,0)$, the vertex $A$).
For $f_{7}$, both $\partial_{u}f_{7}$ and $\partial_{v}f_{7}$ are strictly negative. Evaluating $f_{7}$ along edge $L_{6,7}$ shows that
$f_{7}\Bigl(u,\,1+\tfrac{3}{2}(1-\sqrt{5})u\Bigr)=\frac{(2-\sqrt{5})u+\tfrac{3+\sqrt{5}}{2}}{\tfrac{\sqrt{5}-3}{2}\,u+\tfrac{1+\sqrt{5}}{2}}.$
Its derivative is
$\frac{d}{du}f_{7}\Bigl(u,\,1+\tfrac{3}{2}(1-\sqrt{5})u\Bigr)=\frac{\tfrac{\sqrt{5}-1}{2}}{\bigl(\tfrac{\sqrt{5}-3}{2}\,u+\tfrac{1+\sqrt{5}}{2}\bigr)^{2}}>0.$
so $f_{7}$ is strictly increasing along $L_{6,7}$.
Therefore, the maximum on this segment is attained at the endpoint with largest $u$, namely
$(u,v)=\Bigl(\frac{1+\sqrt{5}}{6},\,0\Bigr).$
Case II: $m=5,6$ (the denominator is in $\{{\bm{g}}_{2},{\bm{g}}_{4},{\bm{g}}_{7}\}$). For $m=5$ or $m=6$, the (m+1)(m\!+\!1)-th largest magnitude must be attained at index $2$, 44, or $7$. On $T^{\prime}$, ${\bm{x}}\cdot{\bm{g}}_{2}\geq 0$ and ${\bm{x}}\cdot{\bm{g}}_{7}\geq 0$, while ${\bm{x}}\cdot{\bm{g}}_{4}<0$, so $|{\bm{x}}\cdot{\bm{g}}_{2}|={\bm{x}}\cdot{\bm{g}}_{2}$, $|{\bm{x}}\cdot{\bm{g}}_{7}|={\bm{x}}\cdot{\bm{g}}_{7}$, and $|{\bm{x}}\cdot{\bm{g}}_{4}|=-({\bm{x}}\cdot{\bm{g}}_{4})$. The switching interfaces are
$L_{2,4}:\$$2(5+\sqrt{5})\,u+(5+7\sqrt{5})\,v=10+2\sqrt{5},$
$L_{2,7}:\$$12\sqrt{5}\,u+(7\sqrt{5}+5)\,v=10+2\sqrt{5}.$
Similar to Case I, the corresponding subregion partition is illustrated in Fig. 5.
[…]
Figure 5: Triangle $T^{\prime}$ in $(u,v)$-coordinates and switching lines $L_{2,4}$ and $L_{2,7}$, partitioning $T^{\prime}$ into three subregions.
(A) Maximizers in $T_{1}$. Maximizing $f_{2}$ and $f_{7}$ in $T_{1}$ is already done in case I, thus we skip them.
For $f_{4}$, From Lemma IV.5, $\partial_{v}f_{4}<0$. Thus the maximizer is on $v=0$, which is edge CACA
$f_{4}(u,0)=\frac{(10-2\sqrt{5})u+(5+2\sqrt{5})}{(5-\sqrt{5})u+\sqrt{5}}.$
Its derivative is
$\frac{d}{du}f_{4}(u,0)=\frac{5(\sqrt{5}-5)}{\bigl((5-\sqrt{5})u+\sqrt{5}\bigr)^{2}}<0.$
Therefore $f_{4}(u,0)$ is strictly decreasing on $u\in[0,1]$, and the maximum on the edge $v=0$ is attained at $u=0$.
(B) Maximizers in $T_{2}$.
For $f_{2}$, both $\partial_{u}f_{2}$ and $\partial_{v}f_{2}$ are strictly positive. Evaluating $f_{2}$ along edge $L_{2,4}$ shows that
$f_{2}\Bigl(u,\frac{10+2\sqrt{5}-2(5+\sqrt{5})\,u}{5+7\sqrt{5}}\Bigr)=\frac{(-10+10\sqrt{5})\,u+(25+11\sqrt{5})}{(-10+2\sqrt{5})\,u+(15+5\sqrt{5})}.$
Its derivative is
$\frac{d}{du}\,f_{2}(u,v(u))=\frac{240+160\sqrt{5}}{\bigl((-10+2\sqrt{5})\,u+(15+5\sqrt{5})\bigr)^{2}}>0.$
Therefore, the maximum is attained at the endpoint with the largest $u$, i.e., at the vertex $A$.
For $f_{4}$, we have $\partial_{v}f_{4}<0$. Hence any maximizer must lie on $CA\cap T_{2}$ or on $L_{2,7}$. From the previous calculation, along $CA$ we also have $\partial_{u}f_{4}<0$. Therefore, if the maximizer lies on $CA\cap T_{2}$, it must occur at the endpoint where $CA\cap T_{2}$ meets $L_{2,7}$. Consequently, it suffices to restrict attention to $L_{2,7}$. Evaluating $f_{4}$ along edge $L_{2,7}$ shows that
$f_{4}\Bigl(u,\frac{10+2\sqrt{5}-12\sqrt{5}\,u}{7\sqrt{5}+5}\Bigr)=\frac{(13-5\sqrt{5})u+(13+6\sqrt{5})}{(5-7\sqrt{5})u+(5+4\sqrt{5})}.$
Its derivative is
$\frac{d}{du}\,f_{4}=\frac{110+88\sqrt{5}}{\bigl((5-7\sqrt{5})u+(5+4\sqrt{5})\bigr)^{2}}>0.$
Therefore the maximum of $f_{4}$ over $L_{2,7}$ is attained at the endpoint where $v=0$, i.e.,
$(u,v)=\Bigl(\frac{1+\sqrt{5}}{6},\,0\Bigr).$
For $f_{7}$, both $\partial_{u}f_{7}$ and $\partial_{v}f_{7}$ are strictly negative. Thus the maximum is on edge $L_{2,7}$, which has already been calculated in Case I.
(C) Maximizers in $T_{3}$.
The maximuzer for $f_{2}$ in $T_{3}$ is on edge $AB$, maximizing $f_{2}$ along $AB$ is already done in previous case.
For $f_{4}$, since $\partial_{v}f_{4}<0$, the mazimizer is on edge $L_{2,4}$. Since on that edge, $f_{2}=f_{4}$, it is equivalent as evaluating $f_{2}$ along $L_{2,4}$, which is already done in previous case.
For $f_{7}$, since $\partial_{v}f_{7}<0$, the mazimizer is on edge $L_{2,4}$.
$f_{7}\Bigl(u,\frac{10+2\sqrt{5}-2(5+\sqrt{5})\,u}{5+7\sqrt{5}}\Bigr)=\frac{(10\sqrt{5}-10)\,u+(25+11\sqrt{5})}{20u+(15+5\sqrt{5})}.$
Differentiating gives
$\frac{d}{du}\,f_{7}=-\frac{40(10+3\sqrt{5})}{\bigl(20u+(15+5\sqrt{5})\bigr)^{2}}<0.$
Therefore, the maximum of $f_{7}$ over $L_{2,4}$ is attained at $u=0$, i.e.,
$(u,v)=\Bigl(0,\ \frac{1+3\sqrt{5}}{11}\Bigr).$
Case III: $m=7$ (the denominator is in $\{{\bm{g}}_{8},{\bm{g}}_{10}\}$).
[…]
Figure 6: Triangle $T^{\prime}$ in $(u,v)$-coordinates, the switching line $L_{8,10}^{-}$, and the zero line $L_{8}:\ x\cdot g_{8}=0$.
The switching condition $|{\bm{x}}\cdot{\bm{g}}_{8}|=|{\bm{x}}\cdot{\bm{g}}_{10}|$ splits into
$L_{8,10}^{+}:\ {\bm{x}}\cdot{\bm{g}}_{8}={\bm{x}}\cdot{\bm{g}}_{10},\qquad L_{8,10}^{-}:\ {\bm{x}}\cdot{\bm{g}}_{8}=-{\bm{x}}\cdot{\bm{g}}_{10},$
which reduce to
$L_{8,10}^{+}:\ v=-2u,\qquad L_{8,10}^{-}:\ v=(2\sqrt{5}-4)(1-u).$
The switching line $L_{8,10}^{+}:\ v=-2u$ intersects $T^{\prime}$ only at the vertex $C=(0,0)$. Hence, it suffices to consider the two subregions partitioned by
$L_{8,10}^{-}:\ v=(2\sqrt{5}-4)(1-u),$
as illustrated in Fig. 6. We further denote by $L_{8}$ the zero line of ${\bm{x}}\cdot{\bm{g}}_{8}$, i.e., ${\bm{x}}\cdot{\bm{g}}_{8}=0$.
In subregion $T_{1}$, the ordering satisfies
$|{\bm{x}}\cdot{\bm{g}}_{10}|\geq|{\bm{x}}\cdot{\bm{g}}_{8}|.$
Moreover, since ${\bm{x}}\cdot{\bm{g}}_{10}\geq 0$ throughout $T^{\prime}$, the $m$-height function for $m=7$ reduces to
$f_{10}(u,v)=\frac{{\bm{x}}\cdot{\bm{g}}_{1}}{{\bm{x}}\cdot{\bm{g}}_{10}}.$
By Lemma IV.5, we have
$\partial_{u}f_{10}<0,\qquad\partial_{v}f_{10}>0\quad\text{on }T^{\prime}.$
Therefore, the maximum of $f_{10}$ over $T_{1}$ is attained at the vertex with minimal $u$ and maximal $v$, namely at the intersection of $L_{8,10}^{-}$ and the edge $BC$, which is
$(u,v)=(0,\ 2\sqrt{5}-4).$
In subregion $T_{2}$, the ordering is reversed and the $m$-height function is
$f_{8}(u,v)=\frac{{\bm{x}}\cdot{\bm{g}}_{1}}{-\,{\bm{x}}\cdot{\bm{g}}_{8}}.$
Then direct differentiation yields
$\partial_{u}f_{8}$ $=\frac{\Bigl(1-\frac{\sqrt{5}}{5}\Bigr)v-(1+\sqrt{5})}{D(u,v)^{2}}<0,$
$\partial_{v}f_{8}$ $=\frac{\Bigl(\frac{\sqrt{5}}{5}-1\Bigr)u-\Bigl(\frac{3}{2}+\frac{7\sqrt{5}}{10}\Bigr)}{D(u,v)^{2}}<0.$
Hence $f_{8}$ is strictly decreasing in both variables, and its maximum over $T_{2}$ must be attained on the boundary $L_{8,10}^{-}$.
Evaluating $f_{8}$ along $L_{8,10}^{-}:\ v=(2\sqrt{5}-4)(1-u)$ gives
$f_{8}\Bigl(u,\ (2\sqrt{5}-4)(1-u)\Bigr)=\frac{(3-\sqrt{5})u+\sqrt{5}}{(3-\sqrt{5})u+(\sqrt{5}-2)}.$
Differentiating,
$\frac{d}{du}\,f_{8}\Bigl(u,\ (2\sqrt{5}-4)(1-u)\Bigr)=\frac{2(\sqrt{5}-3)}{\bigl((3-\sqrt{5})u+(\sqrt{5}-2)\bigr)^{2}}<0,$
so this restriction is strictly decreasing in $u$. Consequently, the maximum of $f_{8}$ on $T_{2}$ is attained at the endpoint $u=0$, i.e.,
$(u,v)=(0,\ 2\sqrt{5}-4).$
The following table summarizes the $m$-height profiles of the dual polyhedral codes.
TABLE I: $m$-height profiles of the dual polyhedral codes.
Code $m$ $m$-height
Dual icosahedral 11 $\sqrt{5}$
22 $\sqrt{5}$
33 $2+\sqrt{5}$
Dual dodecahedral 11 $\dfrac{3}{\sqrt{5}}$
22 $\varphi$
33 $4-\sqrt{5}$
44 33
55 33
66 $2+\dfrac{3\sqrt{5}}{5}$
77 $5+2\sqrt{5}$